# Search for $C$-even states decaying to $D_{s}^{\pm}D_{s}^{*\mp}$ with masses between $4.08$ and $4.32$ $\rm GeV/{\it c}^{2}$

### v1

@@ Abstract
using the MC sample with $\Gamma=50~{}\mathrm{MeV}$ and $J^{PC}=0^{-+}$, and they are fixed in the latter fit to $M_{\gamma_{1}}^{\rm rec}$. The $m_{1}$ and $\sigma_{1}$ dependence on the mass of the $X$ state are shown in Fig. 2. A parameterization of the resolution, as a function of $m_{\rm 1}$ and $\sigma_{\rm 1}$ vs. $M_{\rm X}$ distributions are fitted with a first order polynomial functions. The $m_{\rm 1}$ and $\sigma_{\rm 1}$ dependences for other widths and $J^{PC}$ are determined separately. The trends are similar to those shown in Fig. 2. The mass resolutions of the $X$ in the region $[4.08,~{}4.32]~{}{\rm GeV}/c^{2}$ are determined from these fitted function. The upper limits of the number of signal events as a function of the $X$ mass for different $X$ widths and quantum numbers are shown in Fig. 3.
Refer to captionRefer to caption
Fig. 2: Distributions of $m_{1}$ (upper) and $\sigma_{1}$ (lower) for $\Gamma=50~{}\mathrm{MeV}$ and $J^{PC}$ = $0^{-+}$ according to MC simulation studies. The solid lines are the fit results.
Fig. 3: Upper limits on the number of $X$ signal events as a function of $X$ mass for different quantum numbers ($J^{PC}=0^{-+}$, $1^{-+}$, $1^{++}$, $2^{-+}$, and $2^{++}$) and widths of the $X$ ($\Gamma=5$, $10$, $50$, $80$, and $100~{}\mathrm{MeV}$).
@@
where $N_{\rm sig}$ is the number of signal events obtained from the fit; $\bar{\epsilon}=\sum_{j,k=1}^{2}\epsilon_{jk}\mathcal{B}_{j}\mathcal{B}_{k}$ represents the average detection efficiency, where $\epsilon_{jk}$ is the detection efficiency for $e^{+}e^{-}\to\gamma X,X\to D_{s}^{\pm}D_{s}^{*\mp},D_{s}^{*\mp}\to\gamma D_{s}^{\mp},$ $D_{s}^{+}\to j,D_{s}^{-}\to k$; and $\mathcal{B}_{j}$ ($\mathcal{B}_{k}$) is the branching fraction of $D_{s}^{+}\to j$ ($D_{s}^{-}\to k$). Here $j$ and $k$ are the first and second decay modes of the $D_{s}$. The ISR correction factor ($f^{r}=(1+\delta)$) is calculated following the procedure as described in Refs. ref:kkmca ; ref:kkmcb . The vacuum polarization correction factor is $f^{v}=|1+\Pi|^{2}$ Jegerlehner:2011mw $\mathcal{B}[{D_{s}^{*}\to\gamma D_{s}]}$ is the branching fraction of $D_{s}^{*}\to\gamma D_{s}$. The results are summarized in Table 2.
Table 2: The upper limits of Born cross section of $e^{+}e^{-}\to\gamma X$ times the branching fraction of $X\to D_{s}^{\pm}D_{s}^{*\mp}$ at 90% C.L. for each candidate $C$-even state. Here, $f^{r}$, $f^{v}$, $N_{\rm sig}$, $N_{\rm sig}^{\rm UL}$, and $\bar{\epsilon}$ are the ISR correction factor, the vacuum polarization correction factor, the number of signal events, the upper limit on the number of signal events, and the average detection efficiency, respectively. “$\sigma^{\mathrm{UL}}\cdot\mathcal{B}$ with sys.” stands for the upper limits of the cross section times the branching ratio with systematic uncertainties. “Significance” represents statistical significance.
@@
$\sigma^{\rm UL}\cdot\mathcal{B}$ with sys. (pb) $13.3$ $36.3$ $45.5$ $4.1$ $15.7$ $51.7$
Figure 4 illustrates the upper limits of $\sigma^{\rm UL}\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ under different assumptions of the mass, width, and $J^{PC}$ of the $X$ states after considering all the systematic uncertainties. Detailed descriptions of the systematic uncertainties are provided in the subsequent sections. The detection efficiency as a function of $M$ for different width and $J^{PC}$ assumptions of the $X$ state is determined from Group-II MC simulations. The distributions are fitted with a first order polynomial function, as shown in Fig. 5.
Fig. 4: Upper limits on $\sigma[e^{+}e^{-}\to\gamma X]\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ as a function of $X$ mass for different quantum numbers ($J^{PC}=0^{-+}$, $1^{-+}$, $1^{++}$, $2^{-+}$, and $2^{++}$) and widths of the $X$ ($\Gamma=5$, $10$, $50$, $80$, and $100~{}\mathrm{MeV}$) after considering the systematic uncertainties.
Fig. 5: Distributions of the efficiencies for the four channels for different $J^{PC}$ and width assumptions. The lines are the fit results.
@@
The systematic uncertainties on $\sigma\cdot\mathcal{B}$ are mainly from the luminosity measurement, the detection efficiency, the quoted branching fractions, the ISR correction factor, and the fit models. The uncertainty from the vacuum polarization correction is negligible. A summary of systematic uncertainties is given in Table 3. The total systematic uncertainty is taken as a quadrature sum of each contribution.
@@
Efficiency curve parameterization $0.7$ $1.4$ $1.5$ $1.0$ $0.9$
@@
(b) The difference in the tracking efficiencies between data and MC simulation is estimated to be $1.0\%$ per track Ablikim:092009 ; Ablikim:2011kv ; error_ppb2pi . There are six charged tracks in DC-I, four tracks in DC-II and DC-III, and two tracks in DC-IV. The detection efficiencies of the four channels are changed by ±m%plus-or-$\pm m\%$ to recalculate the Born cross section, where $m$ is the number of charged tracks in the channel. The larger difference to the nominal result, $5.7\%$, is taken as the systematic uncertainty.
(c) The systematic uncertainty from PID is $1.0\%$ per $\pi/K$ track Ablikim:092009 ; Ablikim:2011kv ; error_ppb2pi . This uncertainty is calculated using the same method as used for tracking, and $5.7\%$ is taken as the systematic uncertainty.
(d) The systematic uncertainty of the photon reconstruction efficiency is $1.0\%$ per photon ref:gamma based on studies of the processes $\chi_{0,2}\to\pi^{0}\pi^{0},\eta\eta$. Because there are two photons in the final state, $2.0\%$ is assigned as the systematic uncertainty.
@@
(g) The systematic uncertainties from the $D_{s}$ and $D_{s}^{*}$ mass window requirements are due to the mass resolution difference between data and MC. It is estimated by correcting the mass resolution in MC to improve the consistency between data and MC simulation. The correction is applied by smearing the MC distribution with a Gaussian function, and the parameters of the Gaussian function are obtained from the control sample of $e^{+}e^{-}\to D_{s}^{\pm}D_{s}^{*\mp},~{}D_{s}^{*\mp}\to\gamma D_{s}^{\mp}$. The mean and standard deviation of the Gaussian function are $(0.3\pm 0.6)~{}\mathrm{MeV}/c^{2}$ and $(3.7\pm 0.9)~{}\mathrm{MeV}/c^{2}$ for the $D_{s}$, $(1.5\pm 1.1)~{}\mathrm{MeV}/c^{2}$ and $(0.5\pm 4.2)~{}\mathrm{MeV}/c^{2}$ for the $D_{s}^{*}$. The selection efficiency difference with or without the smearing procedure is taken as the systematic uncertainty, which is $0.3\%$ and $0.1\%$ for the $D_{s}$ and $D_{s}^{*}$, respectively.
@@
(h) In the nominal results, the detection efficiency and the ISR correction factor are obtained by assuming the Born cross section of $e^{+}e^{-}\to\gamma X$ follows the line shape of the $\psi(4660)$, which is a well-established vector state in this energy region. Different shapes of the input cross sections result in different probabilities of generating events with the ISR photon. Assuming the branching fraction of $\psi(4660)\to\gamma X$ is small, the input cross section line shape is changed to $E_{\gamma}^{3}/s$ (similar to the pure continuum process), where $E_{\gamma}$ is energy of the radiative photon and $s$ is the square of the center-of-mass energy. There is a large difference between the two input line shapes, which leads to large differences in the detection efficiencies and the ISR correction factors. The changes in the resultant Born cross sections are taken as the systematic uncertainty from the ISR correction factor ref:weight .
@@
(i) For the calculation of the Born cross section of $e^{+}e^{-}\to\gamma X,X\to D_{s}^{\pm}D_{s}^{*\mp}$ with different $M$ under assumed $\Gamma$ and $J^{PC}$, the detection efficiency is determined from the fitted curves. The uncertainty in the efficiency induced by the fitted parameters is $\Delta\epsilon_{jk}=\sqrt{(m\cdot\Delta a_{jk})^{2}+(\Delta b_{jk})^{2}+2\cdot m\cdot\mathrm{cov}(a_{jk},b_{jk})}$, where $a_{jk}$ and $b_{jk}$ are the parameters of the polynomial function, $\Delta a_{jk}$ and $\Delta b_{jk}$ are the corresponding uncertainties, and $\mathrm{cov}(a_{jk},b_{jk})$ is the covariance between $a_{jk}$ and $b_{jk}$ given by the corresponding fit. The detection efficiency, $\epsilon_{jk}$, is varied by ±$\pm\Delta\epsilon_{jk}$, and the difference on the efficiency is taken as the systematic uncertainty.
@@
(j) The systematic uncertainty from the parameterization of the efficiency curves is estimated by changing the fitted function from a first order to a second order polynomial function. The average detection efficiency is recalculated with parameters read from the alternative fitted curves. The largest difference is taken as the systematic uncertainty.
@@
(k) The branching fractions of $D_{s}\to K^{+}K^{-}\pi$, $D_{s}\to K^{0}_{S}K$, $D_{s}^{*}\to\gamma D_{s}$, and $K^{0}_{S}\to\pi^{+}\pi^{-}$ are taken from the PDG PDG2018 , and their uncertainties are propagated to the Born cross section measurement. The systematic uncertainty caused by the branching fractions of $D_{s}\to K^{+}K^{-}\pi$, $D_{s}\to K^{0}_{S}K$, $D_{s}^{*}\to\gamma D_{s}$, and $K^{0}_{S}\to\pi^{+}\pi^{-}$ are $3.3\%$, $0.8\%$, $0.8\%$, and $0.1\%$, respectively.
@@
(l) The fit-related systematic uncertainties are estimated by varying the fit models. The largest upper limit on the number of signal events is taken as the upper limit of the Born cross section with the systematic uncertainty considered. The line shapes of the background contributions from $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm(*)}D_{s}^{*\mp}$ are determined from the MC simulations with the latest precise measurements as input. The uncertainties are found to be negligible. The uncertainty from the description of the remaining background events is estimated by changing the constant polynomial function to the line shape extracted from generic MC simulation. The value listed under $1^{++}$ is determined from $\chi_{c1}(3P)$. In the determination of the upper limit of the number of signal events with different $M$ under assumed width and $J^{PC}$, the mass resolutions are read from the fit curves shown in Fig. 2. The systematic uncertainties from $m_{1}-m_{2},\sigma_{2}/\sigma_{1}$, and $f$ are considered by varying the parameters within their uncertainties, while the systematic uncertainties from $m_{1}~{}{\rm and}~{}\sigma_{1}$, are considered by varying $\Delta m_{1}$ and $\Delta\sigma_{1}$. The largest upper limit on the number of signal events is taken. Here $\Delta m_{1}$ and $\Delta\sigma_{1}$ are calculated using the same formula as used to calculate $\Delta\epsilon_{jk}$. To take into account the systematic uncertainties in the upper limit calculation, the likelihood distributions are convolved with Gaussian functions as
@@
In summary, based on a data sample collected at $\sqrt{s}=(4681.92\pm 0.08\pm 0.29)~{}\mathrm{MeV}$ with an integrated luminosity of $(1667.39\pm 0.21\pm 8.84)~{}\mathrm{pb}^{-1}$, a search for $C$-even states is performed via the process $e^{+}e^{-}\to\gamma D_{s}^{\pm}D_{s}^{*\mp}$. No statistically significant signals are observed. Upper limits on $\sigma[e^{+}e^{-}\to\gamma X]\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ ($X$= $\eta_{c2}(2D)$$\chi_{c1}(3P)$$\chi_{c2}(3P)$$X(4080)$$X(4217)$, and $X(4279)$) at $90\%$ C.L. are determined to be $13.3$ pb, $36.3$ pb, $45.5$ pb, $4.1$ pb, $15.7$ pb, and $51.7$ pb, respectively. Upper limits with other mass and width assumptions are also determined and shown in Fig. 4. These upper limits can be used to constrain theoretical models.

### v2

@@ Abstract
using the MC sample with $\Gamma=50~{}\mathrm{MeV}$ and $J^{PC}=0^{-+}$, and they are fixed in the latter fit to $M_{\gamma_{1}}^{\rm rec}$. The $m_{1}$ and $\sigma_{1}$ dependence on the mass of the $X$ state are shown in Fig. 2. A parametrization of the resolution, as a function of $m_{1}$ and $\sigma_{1}$ vs. $M_{\rm X}$ distributions, is fitted with first order polynomial functions. The dependencies of $m_{1}$ and $\sigma_{1}$ on other width and $J^{PC}$ are studied separately. The trends are similar to those shown in Fig. 2. The mass resolutions of the $X$ in the region $[4.08,~{}4.32]~{}{\rm GeV}/c^{2}$ are evaluated from these fitted function. The upper limits of the number of signal events as a function of the $X$ mass for different $X$ widths and quantum numbers are shown in Fig. 3.
Fig. 2: $M_{\rm X}$ dependence of $m_{1}$ (upper) and $\sigma_{1}$ (lower) for $\Gamma=50~{}\mathrm{MeV}$ and $J^{PC}$ = $0^{-+}$ according to MC simulation studies. The solid lines are the fit results.
Fig. 3: Upper limits on the number of $X$ signal events as a function of $X$ mass for different quantum numbers ($J^{PC}=0^{-+}$, $1^{-+}$, $1^{++}$, $2^{-+}$, and $2^{++}$) and widths of the $X$ ($\Gamma=5~{}\mathrm{MeV}$, $10~{}\mathrm{MeV}$, $20~{}\mathrm{MeV}$, $50~{}\mathrm{MeV}$, $80~{}\mathrm{MeV}$, and $100~{}\mathrm{MeV}$).
@@
where $N_{\rm sig}$ is the number of signal events obtained from the fit; $\bar{\epsilon}=\sum_{j,k=1}^{2}\epsilon_{jk}\mathcal{B}_{j}\mathcal{B}_{k}$ represents the average detection efficiency, where $\epsilon_{jk}$ is the detection efficiency for $e^{+}e^{-}\to\gamma X,X\to D_{s}^{\pm}D_{s}^{*\mp},D_{s}^{*\mp}\to\gamma D_{s}^{\mp},$ $D_{s}^{+}\to j,D_{s}^{-}\to k$; and $\mathcal{B}_{j}$ ($\mathcal{B}_{k}$) is the branching fraction of $D_{s}^{+}\to j$ ($D_{s}^{-}\to k$). Here $j$ and $k$ are the first and second decay modes of the $D_{s}$. The ISR correction factor ($f^{r}=(1+\delta)$) is calculated following the procedure described in Refs. ref:kkmca ; ref:kkmcb . The vacuum polarization correction factor is $f^{v}=1/|1-\Pi|^{2}$ Jegerlehner:2011mw $\mathcal{B}[{D_{s}^{*}\to\gamma D_{s}]}$ is the branching fraction of $D_{s}^{*}\to\gamma D_{s}$. The results are summarized in Table 2. The product of the upper limit of the cross section and the branching fraction ranges between $4$ pb and $35$ pb, depending on the mass and width of the $X$ states.
Table 2: Product of upper limits of Born cross section of $e^{+}e^{-}\to\gamma X$ and the branching fraction of $X\to D_{s}^{\pm}D_{s}^{*\mp}$ at 90% C.L. for each candidate $C$-even state, where $f^{r}$, $f^{v}$, $N_{\rm sig}$, $N_{\rm sig}^{\rm UL}$, $N_{\mathrm{bkg}1,2,4}$, and $\bar{\epsilon}$ are the ISR correction factor, the vacuum polarization correction factor, the number of signal events, the upper limit on the number of signal events, the numbers of background events, and the average detection efficiency, respectively. “$\sigma^{\mathrm{UL}}\cdot\mathcal{B}$ with sys.” stands for the upper limits of the cross section times the branching ratio with systematic uncertainties. “Significance” represents statistical significance.
@@
$\sigma^{\rm UL}\cdot\mathcal{B}$ with sys. (pb) $10.7$ $23.4$ $28.3$ $4.1$ $13.5$ $32.4$
Figure 4 illustrates the upper limits of $\sigma^{\rm UL}\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ under different assumptions of the mass, width, and $J^{PC}$ of the $X$ states after considering all the systematic uncertainties. Detailed descriptions of the systematic uncertainties are provided in the subsequent sections. The detection efficiency as a function of $M$ for different width and $J^{PC}$ assumptions of the $X$ state is estimated from Group-II MC simulations. The $M_{X}$ dependence of the detection efficiency is fitted with a first order polynomial function, as shown in Fig. 5.
Fig. 4: Upper limits on $\sigma[e^{+}e^{-}\to\gamma X]\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ as a function of $X$ mass for different quantum numbers ($J^{PC}=0^{-+}$, $1^{-+}$, $1^{++}$, $2^{-+}$, and $2^{++}$) and widths of the $X$ ($\Gamma=5~{}\mathrm{MeV}$, $10~{}\mathrm{MeV}$, $20~{}\mathrm{MeV}$, $50~{}\mathrm{MeV}$, $80~{}\mathrm{MeV}$, and $100~{}\mathrm{MeV}$) after considering the systematic uncertainties.
Fig. 5: $M_{\rm X}$ dependence of the efficiencies for the four channels for different $J^{PC}$ and width assumptions. The lines are the fit results.
@@
The systematic uncertainties on $\sigma\cdot\mathcal{B}$ are mainly originated from the luminosity measurement, the detection efficiency, the quoted branching fractions, the ISR correction factor, and the fit models. The uncertainty from the vacuum polarization correction is negligible. A summary of systematic uncertainties is given in Table 3. The total systematic uncertainty is taken as a quadrature sum of each contribution.
@@
Efficiency curve parametrization $0.7$ $1.4$ $1.5$ $1.0$ $0.9$
@@
(b) The difference in the tracking efficiencies between data and MC simulation is estimated to be $1.0\%$ per track for that decay from IP Ablikim:092009 ; Ablikim:2011kv ; error_ppb2pi . There are six charged tracks in DC-I, four tracks in DC-II and DC-III, and two tracks in DC-IV. The charged tracks ($\pi^{+}\pi^{-}$) of the $K_{S}^{0}$ daughters are not counted. The detection efficiencies of the four channels are changed by ±$\pm m\cdot 1.0\%$ to recalculate the Born cross section, where $m$ is the number of charged tracks in the channel. The larger difference to the nominal result, $5.7\%$, is taken as the systematic uncertainty.
(c) The systematic uncertainty from PID is $1.0\%$ per $\pi/K$ track Ablikim:092009 ; Ablikim:2011kv ; error_ppb2pi . This uncertainty is calculated using the same method as used for tracking efficiency, and $5.7\%$ is taken as the systematic uncertainty.
(d) The systematic uncertainty of the photon reconstruction efficiency is $1.0\%$ per photon ref:gamma based on studies of the processes $\chi_{c0,2}\to\pi^{0}\pi^{0},\eta\eta$. Because there are two photons in the final state, $2.0\%$ is assigned as the systematic uncertainty.
@@
(g) The systematic uncertainties from the $D_{s}$ and $D_{s}^{*}$ mass window requirements are due to the mass resolution difference between data and MC. It is estimated by correcting the mass resolution in MC to improve the consistency between data and MC simulation. The correction is applied by smearing the MC distribution with a Gaussian function, and the parameters of the Gaussian function are obtained from the control sample of $e^{+}e^{-}\to D_{s}^{\pm}D_{s}^{*\mp},~{}D_{s}^{*\mp}\to\gamma D_{s}^{\mp}$. The mean and standard deviation of the Gaussian function are $(0.3\pm 0.6)~{}\mathrm{MeV}/c^{2}$ and $(3.7\pm 0.9)~{}\mathrm{MeV}/c^{2}$ for the $D_{s}$, $(1.5\pm 1.1)~{}\mathrm{MeV}/c^{2}$ and $(0.5\pm 4.2)~{}\mathrm{MeV}/c^{2}$ for the $D_{s}^{*}$. The selection efficiency difference with or without (used as nominal result) the smearing procedure is taken as the systematic uncertainty, which is $0.3\%$ and $0.1\%$ for the $D_{s}$ and $D_{s}^{*}$, respectively.
@@
(h) In the nominal results, the detection efficiency and the ISR correction factor are obtained by assuming the Born cross section of $e^{+}e^{-}\to\gamma X$ follows the line shape of the $\psi(4660)$, which is a well-established vector state in this energy region. Different shapes of the input cross sections result in different probabilities of generating events with the ISR photon. Assuming the branching fraction of $\psi(4660)\to\gamma X$ is small, the input cross section line shape is changed to $E_{\gamma}^{3}/s$ ($E_{\gamma}^{3}$ term comes from the radiative transition rate ref:es and $1/s$ is based on the assumption that the matrix element exhibits a similar c.m. energy dependence as the pure continuum process), where $E_{\gamma}$ is energy of the radiative photon and $s$ is the square of the center-of-mass energy. There is a large difference between the two input line shapes, which leads to large differences in the detection efficiencies and the ISR correction factors. The changes in the resultant Born cross sections are taken as the systematic uncertainty from the ISR correction factor ref:weight .
@@
(i) For the calculation of the Born cross section of $e^{+}e^{-}\to\gamma X,X\to D_{s}^{\pm}D_{s}^{*\mp}$ with different $M$ under assumed $\Gamma$ and $J^{PC}$, the detection efficiency is read from the fitted curves. The uncertainty in the efficiency induced by the fitted parameters is $\Delta\epsilon_{jk}=\sqrt{(m\cdot\Delta a_{jk})^{2}+(\Delta b_{jk})^{2}+2\cdot m\cdot\mathrm{cov}(a_{jk},b_{jk})}$, where $a_{jk}$ and $b_{jk}$ are the parameters of the polynomial function, $\Delta a_{jk}$ and $\Delta b_{jk}$ are the corresponding uncertainties, and $\mathrm{cov}(a_{jk},b_{jk})$ is the covariance between $a_{jk}$ and $b_{jk}$ given by the corresponding fit. The detection efficiency, $\epsilon_{jk}$, is varied by ±$\pm\Delta\epsilon_{jk}$, and the difference on the efficiency is taken as the systematic uncertainty.
@@
(j) The systematic uncertainty from the parametrization of the efficiency curves is estimated by changing the fitted function from a first-order to a second-order polynomial function. The average detection efficiency is recalculated with parameters from the alternative fitted curves. The largest difference between the nominal and each fit result is taken as the systematic uncertainty.
@@
(k) The branching fractions of $D_{s}\to K^{+}K^{-}\pi$, $D_{s}\to K^{0}_{S}K$, $D_{s}^{*}\to\gamma D_{s}$, and $K^{0}_{S}\to\pi^{+}\pi^{-}$ are taken from the PDG PDG2018 , and their uncertainties are propagated to the Born cross section measurement. The systematic uncertainties caused by the branching fractions of $D_{s}\to K^{+}K^{-}\pi$, $D_{s}\to K^{0}_{S}K$, $D_{s}^{*}\to\gamma D_{s}$, and $K^{0}_{S}\to\pi^{+}\pi^{-}$ are $3.3\%$, $0.8\%$, $0.8\%$, and $0.1\%$, respectively.
@@
(l)The fit-related systematic uncertainties are estimated by varying the fit models. The largest upper limit on the number of signal events is taken as the upper limit of the Born cross section with the systematic uncertainty considered. The line shapes of the background contributions from $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{\pm(*)}D_{s}^{*\mp}$ are taken from the MC simulations with the latest precise measurements as input. The uncertainties are found to be negligible. The systematic uncertainty from the fixed number of background events for the processes $e^{+}e^{-}\to(\gamma_{\rm ISR})D_{s}^{(*)}D_{sJ}$ is estimated by modifying $N_{\mathrm{bkg}3}$ with ±$\pm 1\cdot\sigma$.  The uncertainty arising from the modeling of the remaining background events is accessed by replacing the zero-order polynomial function with the line shape extracted from generic MC simulation. In the determination of the upper limit of the number of signal events with different $M$ under assumed width and $J^{PC}$, the mass resolutions are read from the fit curves shown in Fig. 2. The systematic uncertainties from $m_{1}-m_{2},\sigma_{2}/\sigma_{1}$, and $f$ are considered by varying the parameters within their uncertainties, while the systematic uncertainties from $m_{1}~{}{\rm and}~{}\sigma_{1}$, are considered by varying $\Delta m_{1}$ and $\Delta\sigma_{1}$. The largest upper limit on the number of signal events is taken, where $\Delta m_{1}$ and $\Delta\sigma_{1}$ are calculated using the same formula as used to calculate $\Delta\epsilon_{jk}$. To take into account the systematic uncertainties in the upper limit calculation, the likelihood distributions are convolved with Gaussian functions as
@@
In summary, based on a data sample collected at $\sqrt{s}=(4681.92\pm 0.08\pm 0.29)~{}\mathrm{MeV}$ with an integrated luminosity of $(1667.39\pm 0.21\pm 8.84)~{}\mathrm{pb}^{-1}$, a search for $C$-even states is performed via the process $e^{+}e^{-}\to\gamma D_{s}^{\pm}D_{s}^{*\mp}$. No statistically significant signals are observed. Upper limits on $\sigma[e^{+}e^{-}\to\gamma X]\cdot\mathcal{B}[X\to D_{s}^{\pm}D_{s}^{*\mp}]$ ($X$= $\eta_{c2}(2D)$$\chi_{c1}(3P)$$\chi_{c2}(3P)$$X(4080)$$X(4217)$, and $X(4279)$) at $90\%$ C.L. are determined to be $10.7$ pb, $23.4$ pb, $28.3$ pb, $4.1$ pb, $13.5$ pb, and $32.4$ pb, respectively. Upper limits with other mass and width assumptions are also calculated and shown in Fig. 4. Furthermore, the product of the upper limit of the cross section and the branching fraction is between $4$ pb to $35$ pb. Due to statistical limitations, the upper limit for $\psi(4660)\to\gamma\chi_{c1,2}(3P)$ is below the sensitivity of the theoretical prediction in Ref. ref:em . There is no other theoretical prediction for the cross sections of other $C$-even states. More data samples and theoretical calculations are needed to investigate the EM transition properties of $\psi(4660)$.